# GVE-LPA: Fast Label Propagation Algorithm (LPA) for Community Detection in Shared Memory Setting

### v4

@@ 1. Introduction
Community detection is the problem of identifying cohesive groups of vertices, also known as clusters, in complex networks that minimize cuts and maximize internal density (Garza and Schaeffer, 2019). This NP-hard problem finds applications in various domains, including drug discovery, protein annotation, topic exploration, anomaly detection, and criminal identification. Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). The Label Propagation Algorithm (LPA), also known as RAK, (Blondel et al., 2008) is a popular heuristic-based approach for community detection, with the modularity metric (Newman, 2006) being used to measure the quality of communities identified.
In the past few years, there has been an unprecedented surge in the gathering of data and the depiction of their interconnections through graphs. This surge has underscored the need for devising efficient parallel algorithms tailored for identifying communities within massive networks. The significance of the multicore/shared memory environment in this context is paramount, owing to its energy-efficient nature and the prevalence of hardware equipped with substantial DRAM capacities. Existing studies on LPA propose algorithmic optimizations and parallelization techniques, but do not study efficient data structures for picking the most weighted label, to the best of our knowledge. Moreover, the proposed techniques are scattered over a number of papers.
1.1. Our Contributions
This report introduces GVE-LPA [^1] https://github.com/puzzlef/rak-communities-openmp, an optimized parallel implementation of LPA for shared memory multicores. On a machine with two 16-core Intel Xeon Gold 6226R processors, GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively. GVE-LPA is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, and achieves a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. With doubling of threads, GVE-LPA exhibits an average performance scaling of 1.7×1.7$\times$1.7 ×.
@@ 2. Related work
The Label Propagation Algorithm (LPA) is a diffusion-based technique for identifying communities. It is faster and more scalable than Louvain, as it does not require repeated optimization steps and is easier to parallelize (Newman, 2004; Raghavan et al., 2007). For parallelization of LPA, vertex assignment has been achieved with guided scheduling (Staudt and Meyerhenke, 2015), parallel bitonic sort (Soman and Narang, 2011), and pre-partitioning of the graph (Kuzmin et al., 2015).
In terms of quality, LPA may generate communities with low modularity, leading to the emergence of a dominant community that engulfs the majority of nodes, obstructing finer community structures (Gregory, 2010). To address instability of result, community structure quality, and performance due to vertex processing order, specific ordering strategies have been attempted. These include updating only active nodes (Xie and Szymanski, 2011) or prioritizing certain subsets based on properties like being core or boundary nodes (Gui et al., 2018). In some cases, oscillations in label assignments may occur instead of convergence. To address this, asynchronous mode (Leung et al., 2009), alternate updates of independent node subsets (Cordasco and Gargano, 2012), and parallel graph coloring techniques (Cordasco and Gargano, 2012) have been attempted. Raghavan et al. (Raghavan et al., 2007) report that after only five iterations, labels of $95\%$ of nodes converge to their final values.
Additional improvements upon the LPA include using a stable (non-random) mechanism of label choosing in the case of multiple best labels (Xing et al., 2014), addressing the issue of monster communities (Berahmand and Bouyer, 2018; Sattari and Zamanifar, 2018), identifying central nodes and combining communities for improved modularity (You et al., 2020), and using frontiers with alternating push-pull to reduce the number of edges visited and improve solution quality (Liu et al., 2020). A number of variants of LPA have been proposed, but the original formulation is still the simplest and most efficient (Garza and Schaeffer, 2019).
@@ 4.1. Optimizations for LPA
Our optimizations include using OpenMP’s dynamic loop schedule, setting an initial tolerance of $0.05$, enabling vertex pruning, employing the strict version of LPA, and using fast collision-free per-thread hashtables which are well separated in their memory addresses (Far-KV). See below for the details on each optimization.
We evaluate multiple alternatives for each optimization, and show the relative time and the relative modularity of communities obtained by each alternative in Figure 3. We perform these tests on every graph in the dataset (refer to Table 1), conducting them five times on each graph to minimize the influence of noise. We then calculate the geometric mean for the runtime and arithmetic mean for the modularity, and represent them as ratios within each optimization category.
@@ 5.1. Experimental Setup
Table 1. List of $13$ graphs obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with ***). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities identified with GVE-LPA.
@@ 5.2. Comparing Performance of GVE-LPA
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. igraph LPA fails to run on protein k-mer graphs, i.e., kmer_A2a and kmer_V1r, and hence its results on these graphs are not shown. GVE-LPA exhibits an average speedup of 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA respectively. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $6.6\%$ / $0.2\%$ higher modularity than FLPA (especially on road networks) and igraph LPA respectively, and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs).
@@ 6. Conclusion
In summary, this study focuses on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. We consider 6 different optimizations, which significantly improve the performance of the algorithm. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively, while identifying communities of $6.6\%$ / $0.2\%$ higher quality than FLPA and igraph LPA respectively, and of $4.1\%$ lower quality than NetworKit LPA.
On a web graph with $3.8$ billion edges, GVE-LPA identifies communities in $2.7$ seconds, and thus achieves a processing rate of $1.4$ billion edges/s. GVE-LPA is ideal for applications requiring a high-speed clustering algorithm while accepting a compromise in clustering quality, as it is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, has a strong scaling factor of 1.7×1.7$\times$1.7 × for every doubling of threads, but identifies communities with on average $10.9\%$ lower modularity than GVE-Louvain. Future research could explore dynamic algorithms for LPA to accommodate evolving graphs in real-world scenarios. This would allow interactive updation of community memberships of vertices on large graphs.

### v5

@@ 1. Introduction
Community detection is the problem of identifying cohesive groups of vertices, also known as clusters, in complex networks that minimize cuts and maximize internal density (Garza and Schaeffer, 2019). This problem finds applications in various domains, including drug discovery, protein annotation, topic exploration, anomaly detection, and criminal identification. Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). One of the difficulties in the community detection problem is the lack of apriori knowledge on the number and size distribution of communities (Blondel et al., 2008).
The Label Propagation Algorithm (LPA), also known as RAK, (Raghavan et al., 2007) is a popular diffusion-based approach for identifying communities. It is faster and more scalable than the Louvain method (Blondel et al., 2008), another high-quality community detection algorithm, as it does not require repeated optimization steps and is easier to parallelize (Newman, 2004; Raghavan et al., 2007). Due to this, LPA has been used in a number of other graph problems and applications, such as, finding connected components (Stergiou et al., 2018), graph partitioning (Slota et al., 2014; Wang et al., 2014; Meyerhenke et al., 2014, 2016, 2017; Bae et al., 2020; Akhremtsev et al., 2020; Slota et al., 2020; Zhang et al., 2020a), hypergraph partitioning (Henne, 2015; Gottesbüren et al., 2021), graph coarsening (Valejo et al., 2020), vertex reordering and graph compression (Boldi et al., 2011), unsupervised part-of-speech tagging (Das and Petrov, 2011), sectionalizing power systems (Aziz et al., 2023). Further, Peng et al. (Peng et al., 2014) observe that, while LPA achieves lower modularity scores (a metric for assessing community quality (Newman, 2004)) than the Louvain method, it attains the highest NMI score when compared to the ground truth. A large number of LPA variants have also been proposed (Li et al., 2015a; Farnadi et al., 2015; Li et al., 2015b; SHEN and YANG, 2016; Mohan et al., 2017; Zhang et al., 2017; Berahmand and Bouyer, 2018a; Ma et al., 2018; Sattari and Zamanifar, 2018b; Zheng et al., 2018; Xu et al., 2019; Zarei et al., 2020; Maleki et al., 2020; Zhang et al., 2020b; El Kouni et al., 2021; Roghani et al., 2021; Zhang et al., 2023). Existing studies on LPA propose a number of algorithmic optimizations and parallelization techniques, but there continue to be a number of issue with each implementation. Further, the race for AI development has caused the prices of GPUs to skyrocket. These issues, and the large applicability of LPA, motivates us to present GVE-LPA, [^1] https://github.com/puzzlef/rak-communities-openmp an efficient implementation of LPA for multicore architectures.
@@ 2. Related work
For parallelization of LPA, vertex assignment has been achieved with guided scheduling (Staudt and Meyerhenke, 2015), parallel bitonic sort (Soman and Narang, 2011), and pre-partitioning of the graph (Kuzmin et al., 2015). Additional improvements upon the LPA include using a stable (non-random) mechanism of label choosing in the case of multiple best labels (Xing et al., 2014), addressing the issue of monster communities (Berahmand and Bouyer, 2018b; Sattari and Zamanifar, 2018a), identifying central nodes and combining communities for improved modularity (You et al., 2020), and using frontiers with alternating push-pull to reduce the number of edges visited and improve solution quality (Liu et al., 2020).
@@
We now discuss the implementation of LPA in NetworKit, igraph, and FLPA. The parallel implementation of LPA in NetworKit is given in NetworKit::PLP::run(). This function starts with a unique label for each node. To check for convergence, they use a tolerance of $10^{-5}$, i.e., the algorithm converges once the labels of less than $0.001\%$ of vertices change. This is also referred to as threshold heuristic. To track active nodes, a boolean flag vector is employed, serving as a vertex pruning optimization. A parallel for loop, which uses OpenMP’s guided schedule, then processes only the active nodes. For storing label weights, an std::map is used for each vertex. In contarst, the LPA implementation of igraph and FLPA is sequential. igraph’s LPA implementation, found in the function igraph_community_label_propagation(), shuffles the node order in each iteration. It keeps track of the dominant labels for each new neighbor of a vertex and selects one of the dominant labels using a random number generator. The hashtable is cleared by iterating through the neighbors. Unlike NetworKit, igraph does not use the pruning optimization. It alternates between label updating and control iterations, checking if the current label of a node is not dominant during control iterations, with no tolerance for non-dominant labels. The FLPA implementation in igraph, detailed in the function igraph_i_community_label_propagation(), uses a dequeue for managing a set of unprocessed vertices, which indicates convergence when empty. FLPA does not shuffle for random node order. If a label change occurs, FLPA considers neighbors that are not in the same community.
However, we identify a few issues with the LPA implementations of NetworKit, igraph, and FLPA. In order to assign a unique label to each node in the graph, NetworKit uses OpenMP’s plain parallel for, i.e., it uses static loop scheduling with a chunk size of $1$. This can result in false sharing as threads make writes to consecutive locations in memory. Next, to keep track of the weights associated with each label for a given vertex, they use an std::map. Our experiments show that this is quite inefficient. We, in contrast, use a per-thread keys list and a per-thread full-size values array, as our hashtable. In order to check for convergence, NetworKit uses a tolerance of $10^{-5}$. However, we observe that a tolerance of $10^{-2}$ is generally obtains community of nearly the same quality (in terms of modularity), but converges much faster. Further, NetworKit uses atomic operations on a shared variable to keep a count of the number of updated vertices. This can introduce contention. We, instead, make use of parallel reduce for this. Finally, NetworKit uses a boolean flag vector to keep track of keep track of the set of active nodes. However, our experiments show that using an 8-bit integer flag vector is more efficient, despite its larger memory footprint.
We now discuss issues in the LPA implementation of igraph. igraph randomly shuffles the processing order of nodes in each iteration. This can get quite expensive. igraph checks for convergence in alternate iterations, and algorithm is considered to have converged only if the labels of all nodes are dominant, i.e., are the most popular. Thus it can take a large number of iterations to converge, with minimal gain in community quality. In addition, igraph does not use vertex pruning optimization, i.e., it does not maintain the set of active nodes. Finally, given multiple dominant labels, igraph select a random dominant label as the label of the given node — however, random number generation is slow. We instead use the dominant label with the smallest ID as the label of the given vertex.
Finally, we discuss the issues in FLPA’s implementation. Similar to igraph LPA, given multiple dominant labels, FLPA select a random dominant label as the label of the given node — however, random number generation is slow. Further, FLPA considers the algorithm to have converged only when there are no active nodes. Thus it can take a large number of iterations to converge, with minimal gain in community quality.
@@ 4.1. Optimizations for LPA
We now evaluate multiple alternatives for a few optimizations, and show the relative time and the relative modularity of communities obtained by each alternative in Figure 3. We perform these tests on every graph in the dataset (refer to Table 1), conducting them five times on each graph to minimize the influence of noise. We then calculate the geometric mean for the runtime and arithmetic mean for the modularity, and represent them as ratios within each optimization category. While none of our optimizations are particularly novel, we carefully address each one and combine them to achieve superior performance compared to other implementations of LPA. We believe our implementation will aid the community in advancing more rapidly and tackling more challenges.
@@ 5.1. Experimental Setup
Table 1. List of $13$ graphs obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with *). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities identified with GVE-LPA.
@@ 5.2. Comparing Performance of GVE-LPA
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. igraph LPA fails to run on protein k-mer graphs, i.e., kmer_A2a and kmer_V1r, and hence its results on these graphs are not shown. GVE-LPA exhibits an average speedup of 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA respectively. While FLPA and igraph LPA are sequential implementations, GVE-LPA, running on $64$ threads, is more than 64×64$\times$64 × faster than these implementations. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $6.6\%$ / $0.2\%$ higher modularity than FLPA (especially on road networks) and igraph LPA respectively, and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs).
@@ 6. Conclusion
In summary, this study focused on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively, while identifying communities of $6.6\%$ / $0.2\%$ higher quality than FLPA and igraph LPA respectively, and of $4.1\%$ lower quality than NetworKit LPA. On a web graph with $3.8$ billion edges, GVE-LPA identifies communities in $2.7$ seconds, and thus achieves a processing rate of $1.4$ billion edges/s. GVE-LPA is ideal for applications requiring a high-speed clustering algorithm while accepting a compromise in clustering quality, as it is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, has a strong scaling factor of 1.7×1.7$\times$1.7 × for every doubling of threads, but identifies communities with on average $10.9\%$ lower modularity than GVE-Louvain.